# PiFlow: Principle-aware Scientific Discovery with Multi-Agent Collaboration

### v2

@@ 4.4 Evaluation Metrics
Two metrics are employed to evaluate the performance of PiFlow, as detailed below:
Solution Quality (SQ). We measure the optimal objective value with a percentage relative to the theoretical maximum value $\mu_{\text{absolute}}$, denoted as:
$\mathrm{SQ}=\frac{\max\{y_{k}|\left<p_{k},y_{k}\right>\in\mathcal{T}_{t}\}}{\mu_{\text{absolute}}}\times 100\%.$ (2)
Specifically, for NHO, the theoretical maximum is reported as $\mu_{\text{absolute}}^{\text{g-factor}}=2.0$ following Greenfield et al. (2021). The larger the g-factor, the stronger the chirality. For MBO, a strict threshold of $\mu_{\text{absolute}}^{\text{pchembl}}=6.5$ has been reported (Lenselink et al., 2017) for lead compound optimization stage, larger value indicates a higher bio-activity and vice versa. For SPO, the reference value is set to be $25^{\circ}\mathrm{C}$ , which is equal to room temperature 298.15K, denoted as $\mu_{\text{absolute}}^{\text{Tc}}=298.15K$.
Area Under the Curve (AUC). Exploration efficiency requires both (1) rapid convergence and (2) high objective values. We quantify these two factors by defining the AUC metric. Given the trajectory $\left<p_{k},y_{k}\right>\in\mathcal{T}_{t}$ across $t$ steps, AUC is computed via the trapezoidal rule. For meaningful comparisons, we normalize by the maximum possible area:
$\mathrm{AUC}=\frac{\sum_{i=1}^{t-1}\frac{y_{i}+y_{i+1}}{2}}{\mu_{\text{absolute}}\cdot(t-1)}\times 100\%$ (3)
In summary, SQ measures the quality of the final outcome, while AUC evaluates the entire discovery process by rewarding both speed and consistency. The detailed rationale for these metric designs is provided in Appendix R.2.
@@ 5.1 Performance Comparison
We use SQ to compare the overall capability of reaching the objective solution, and AUC to assess the efficiency, reflecting progress towards better outcomes over the exploration process.
Table 1 demonstrates that PiFlow achieves significant improvements over all baselines across three benchmarks of NHO, MBO, and SPO. In terms of achieving the target property, measured by SQ, PiFlow consistently leads, outperforming ReAct, MPO, and Vanilla systems by an average of approximately 207.6%, 34.1%, and 94.1%. Additionally, as shown in Figure 3, PiFlow is always the fastest one in reaching the best solution while exploring. More details about models’ response analysis can be seen at Appendix I.
It is worth noting the considerable std observed in our results. This variance is inherent to the stochastic nature of LLM-based agents and can be attributed to the phenomenon of cumulative error. An early suboptimal decision or flawed hypothesis can set an agent on an inefficient discovery trajectory, with reasoning errors compounding over subsequent iterations. This instability highlights why relying solely on the final SQ can be insufficient. In contrast, the AUC provides a more holistic performance measure. A high AUC, as consistently demonstrated by PiFlow, indicates that the agent not only found a high-quality solution but also maintained a robust and efficient path, thereby successfully mitigating the primary risk of cumulative error.
Takeaway: PiFlow substantially outperforms all baselines in SQ and exploration efficiency (AUC). Its strategic robustness is evidenced by its high AUC, which mitigates the cumulative errors of LLMs, and by its success from modest starting points that simulate realistic initial uncertainty (see Appendix P.2).
Table 1: Comparisons between PiFlow and baselines.
Method NHO (g-factor) MBO (pChEMBL) SPO ($T_{c}$)
AUC (%) SQ (%) AUC (%) SQ (%) AUC(%) SQ (%)
ReAct 35.85 $\pm$7.12 41.96 $\pm$0.82 29.61 $\pm$9.74 43.11 $\pm$9.98 5.29 $\pm$0.69 6.41 $\pm$0.96
Vanilla 35.96 $\pm$22.38 46.76 $\pm$7.29 29.71 $\pm$12.66 49.22 $\pm$8.30 11.39 $\pm$11.33 14.16 $\pm$13.37
MPO 43.99 $\pm$2.79 51.29 $\pm$7.77 31.18 $\pm$9.12 57.28 $\pm$5.85 12.68 $\pm$7.75 33.20 $\pm$23.75
PiFlow (ours) 63.51 $\pm$11.18 76.82 $\pm$4.54 46.11 $\pm$16.25 84.55 $\pm$29.63 21.51 $\pm$2.80 34.85 $\pm$1.19
Figure 3: Trajectory comparisons for different optimization methods.
@@ 5.2 Ablation Study
We conduct several ablations to evaluate PiFlow. For these studies, conducted on the NHO task, performance is evaluated based on the same metrics, AUC (%) and SQ (%).
Plug-and-Play. To isolate the direct benefit of PiFlow, we compared the performance with two different LLMs, GPT4.1-mini and Qwen3-32B under the setting of w/ and w/o PiFlow. This is achieved by only including or excluding the steered principle to the Planner Agent via prompt, which then directs subsequent hypothesizing and validation.
As shown in Table 2, for GPT4.1-mini, integrating PiFlow increases the AUC from 37.12% to 41.68% and substantially boosted the SQ from 40.14% to 66.38%. This represents an approximate 12.3% improvement in AUC and a significant 65.4% improvement in SQ. Similarly, for the Qwen3-32B model, the inclusion of PiFlow improves AUC from 27.04% to 37.51% (a 38.7% increase) and SQ from 54.84% to 58.76% (a 7.1% increase).
Table 2: Ablation Study with/without PiFlow
@@
Thought Mode Effect. We also conduct ablations on the internal thought mode of the LLM (referred to as Think in Table 3). This is for agents based on Qwen3-32B and Qwen3-8B models, which support ON/OFF <think>...</think> generation with system prompt including or excluding /no_think. This mode is intended to enable more explicit reasoning steps.
Table 3: Effect of Thinking.
Method/Setting AUC (%) SQ (%)
Qwen3-32B w/ Think 37.51 $\pm$7.70 58.76 $\pm$6.18
w/o Think 45.51 $\pm$11.19 68.86 $\pm$17.67
Qwen3-8B w/ Think 30.55 $\pm$27.45 54.49 $\pm$22.25
w/o Think 42.09 $\pm$16.55 61.59 $\pm$16.43
Interestingly, for both Qwen3-32B and Qwen3-8B, disabling the Thought Mode leads to improved performance, as shown at Table 3. We hypothesize that this phenomenon is due to cognitive fixation. As the key issue is, how do LLM propose scientific hypothesis, forcing the LLM to generate an explicit Chain-of-Thought may cause it to commit prematurely to its own initial line of reasoning. If the first step in its logic is flawed, the entire chain can be led astray, creating a cognitive fixation that is hard to escape. In contrast, disabling the think mode may force the model to rely more on its powerful, holistic pattern-matching capabilities, allowing it to make more intuitive leaps directly from the data (PiFlow’s guidance and the experimental history) to a hypothesis, bypassing potentially flawed intermediate reasoning steps.
Takeaway: Our ablations confirm PiFlow is a robust, plug-and-play enhancement that consistently boosts performance across models. Concurrently, we find that an agent’s internal reasoning is critical: disabling explicit "Think" modes paradoxically improves performance, suggesting that avoiding cognitive fixation allows the LLM to better leverage holistic pattern-matching. This highlights the synergy between high-level strategic guidance from PiFlow and internal reasoning.
5.3 Further Analyses of Robustness
We conducted extensive analyses to probe the robustness and practical utility of PiFlow. The key findings are as follows:
Recovery from poor initialization. The system demonstrates resilience by successfully recovering from deliberately incorrect initial principles, eventually matching the performance of a run guided by correct initialization. This underscores the robustness of the Min-Max (Appendix P).
Temporal dynamics of principle evaluation. PiFlow continuously re-evaluates the utility of scientific principles as new evidence is gathered. This allows it to dynamically discard initially promising but ultimately flawed avenues while elevating principles that prove more fruitful later, showcasing an adaptive balance between exploration and exploitation (Appendix N).
Superior performance against numerical search. On complex tasks with ill-defined search spaces, PiFlow substantially outperforms Bayesian Optimization, achieving an SQ of 84.55% versus BO’s 38.15% in the MBO task, without requiring manual parameter space engineering (Appendix J).
Seamless plug-and-play integration. We validate the practicality of PiFlow through a successful integration with ChemToolAgent (Yu et al., 2024), guiding it to a high-value molecule (pChEMBL of 5.90) without requiring any architectural modifications (Appendix K).
Manageable computational complexity. The core decision mechanism of PiFlow has a computational complexity of $\mathcal{O}(t^{2}\cdot d)$, where $t=|\mathcal{T}_{t}|$ and $d$ is the embedding dimension. This could be optimized to near-linear time for large-scale tasks (Appendix L).
Cost-effectiveness. Superior performance is achieved cost-effectively, reducing token consumption by up to 27% compared to the Vanilla Agent. PiFlow constitutes only 1.5% tokens (Appendix M).
Generalizability and controllability. PiFlow is compatible with various LLMs, and its behavior can be tuned via the $\lambda$ hyperparameter, for which we provide a clear heuristic (Appendices O, Q).
5.4 Theoretical Alignment
To empirically validate the theoretical guarantees (Theorem 1) of PiFlow, we analyze key aspects of its exploration dynamics, including (1) the bound of average regret and (2) the relationship between regret and information gain, as shown in Figure 4.
(Theoretical Prediction 1) Average regret decay with $\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)$. Figure 4a presents the average regret as a function of iterations on a log-log scale. The alignment evidenced by average regret trajectories adhering to the $T^{-0.5}$ decay (fitted $c\cdot T^{-0.5}$ with $c=1.37,r^{2}=0.96$), a pattern most runs consistently follow. Notably, one run (ID=1), after an initial sharp regret decrease, shows a transient increase before resuming decay. This illustrates PiFlow’s robust exploration avoiding potential local optima with a characteristic of its Min-Max strategy.
(Theoretical Prediction 2) As information gain decreases, the expected regret also decreases. As shown in Figure 4b, the scatter plot of regret versus information gain (points colored by iteration) reveals a clear positive association, confirmed by a fitted trend line: lower information gain generally corresponds to lower regret. Early iterations (higher information gain and regret) transition to later iterations (lower information gain and regret).
@@ 5.4 Theoretical Alignment
Figure 4: Empirical Validation of PiFlow’s Theoretical Alignment. (a) Average regret over iterations aligns with the theoretical sublinear decay; (b) Scatter plot with positive trend of observed regret versus information gain, supporting theoretical expectations; (c) An exemplary PiFlow exploration trajectory in NHO task.
Empirical validation of exploration strategy. Figure 4c displays the trajectory of PiFlow on the NHO objective landscape, which is visualized via Principal Component Analysis (PCA) with contours indicating g-factor values. The path demonstrates a principled strategy: initial broad exploration (iterations 1-16), followed by navigating a low-quality "valley" to escape a local optimum (iterations 16-21), and culminating in efficient convergence to a high-g-factor region (iterations 21-24). This progression empirically validates how our theoretical design translates into effective discovery.
Takeaway: Our empirical analysis corroborates the theoretical guarantees of PiFlow, demonstrating that its strategy is both principled and effective. The alignment of regret dynamics with the predicted sublinear rate and its coupling with information gain validates a robust mechanism that translates theoretical efficiency into a practical ability to escape local optima and converge reliably.
@@ 6 Conclusion
In conclusion, we propose PiFlow, a Plug-and-Play module to strategically guide the Hypothesis-Validation loop through a steering mechanism to address challenges of aimless hypothesizing and unclear connections between hypotheses and evidence. Our approach utilizes a Min-Max optimization that explicitly balances exploitation of high-potential principles with exploration of novel hypotheses, guaranteed by a sublinear average regret bound. Extensive experiments demonstrate that PiFlow can adaptively navigate complex hypothesis spaces without premature convergence on suboptimal solutions, yielding significant improvements over baselines. A detailed discussion of its current limitations and future potential is provided in Appendix C. We hope PiFlow can contribute to advancing automated scientific discovery, inspiring further exploration and innovation.
Appendix A Use of Large Language Models
We utilized a large language model to assist with proofreading and polishing the language in this manuscript.
Appendix B Broader Impact
PiFlow addresses critical bottlenecks in AI-driven scientific discovery where uncertainty leads to aimless exploration. Our method innovatively frames discovery as a structured uncertainty reduction problem. By using an information-theoretic approach within a hypothesis-validation loop, PiFlow systematically filters for instructive scientific principles. Ultimately, PiFlow establishes a new paradigm for automated research, enabling more targeted exploration and accelerating the generation of impactful scientific insights. In Materials Discovery, it speeds the identification of novel compounds like advanced nanomaterials or superconductors, as shown in our tasks. For Biological Discovery, it enhances the search for effective molecules and the understanding of complex systems. Its principles promise similar advancements in other data-intensive fields, from chemistry to medical sciences, facing vast and uncertain hypothesis spaces.
Appendix C Limitation and Future Work
While PiFlow shows notable improvements through its principled Min-Max optimization, its practical implementation approximates a key theoretical component. This means the current system may not fully capture all nuances of true, model-based information gain, especially the direct adversarial interplay with all possible manifestations of the unknown evaluation function $f^{*}$ from the theoretical objective.
Future research could explore more direct estimations of mutual information for this heuristic within the PiFlow framework to potentially further enhance its strategic guidance. Furthermore, we observed that disabling the LLM’s “Thought Mode” surprisingly improves performance, suggesting that forced Chain-of-Thought can induce cognitive fixation. This finding motivates the development of more flexible reasoning frameworks for agents, aiming to better balance deliberate logic with intuitive generation.
Appendix D Detailed Methodological Review
@@
A hallmark and fundamental limitation of many contemporary LLM-based agent systems in scientific discovery is their limited generalizability. As detailed in Table 4, these frameworks are often characterized by a tight coupling between their core logic and a specific scientific domain. This means their implementations, tool integrations, and most critically, their prompt engineering strategies are meticulously tailored for a single scenario, such as organic chemistry or materials science. Although systems like The AI Scientist (Lu et al., 2024) and Agent Laboratory (Schmidgall et al., 2025) demonstrate strong capabilities in scientific research, they are specifically designed for AI domain, along with a focus of the whole workflow rather than strategic decision-making problem. Consequently, transferring these systems to a new domain necessitates significant re-engineering, restricting their out-of-the-box applicability and hindering the development of truly universal systems.
Table 4: Comparison of Generalization Limitations in LLM-Agent Systems
@@
(Ghafarollahi & Buehler, 2025) Automating alloy design and discovery using physics-aware multimodal multi-agent AI Alloy design and discovery
@@
MOOSE-Chem (Yang et al., 2024c) Utilizing Large Language Models for rediscovering unseen chemistry scientific hypotheses. Chemistry hypothesis discovery
@@
ProtAgents (Ghafarollahi & Buehler, 2024b) Protein discovery via large language model multi-agent collaborations combining physics and machine learning Protein discovery
@@
Appendix E Distinction from Advanced Prompt Engineering
We define a scientific principle following Definition 3.1. Formally, each principle is represented as a structured text proposition that can be algorithmically scored by our MinMax optimization (as exampled in Figure 1). Unlike arbitrary text strings, these principles is supposed to be with logical consistency, and this structural requirement is what distinguishes them from general prompts. In short, principles represent foundational concepts or established patterns within a domain (e.g., “higher hydrophobicity often correlates with better cell membrane penetration”). These “principles”, whether has been validated or not, can be proposed by experts or, crucially, extracted from the LLM’s own vast pre-trained knowledge. They serve as high-level starting points to generate specific, testable hypotheses.

### v3

@@ 4.4 Evaluation Metrics
To rigorously evaluate PiFlow, we employ two complementary metrics designed to capture both peak performance (effectiveness) and trajectory dynamics (efficiency):
Solution Quality (SQ). We measure the optimal objective value with a percentage relative to the theoretical maximum or reference value $\mu_{\text{ref}}$, denoted as:
$\mathrm{SQ}=\frac{\max\{y_{k}|\left<p_{k},y_{k}\right>\in\mathcal{T}_{t}\}}{\mu_{\text{ref}}}\times 100\%.$
Specifically, for NHO, the theoretical maximum is reported as $\mu_{\text{ref}}^{\text{g-factor}}=2.0$ following Greenfield et al. (2021). The larger the g-factor, the stronger the chirality. For MBO, a strict threshold of $\mu_{\text{ref}}^{\text{pchembl}}=6.5$ has been reported (Lenselink et al., 2017) for lead compound optimization stage in molecular biology, larger value indicates a higher bio-activity and vice versa. For SPO, the reference value is set to be room temperature of 298.15K, denoted as $\mu_{\text{ref}}^{\text{Tc}}=298.15K$.
Area Under the Curve (AUC). Exploration efficiency requires both (1) rapid convergence and (2) high objective values (exploitation). We quantify these two factors by defining the AUC metric. Given the trajectory $\left<p_{k},y_{k}\right>\in\mathcal{T}_{t}$ across $t$ steps, AUC is computed via the trapezoidal rule. For meaningful comparisons, we normalize by the maximum possible area:
$\mathrm{AUC}=\frac{\sum_{i=1}^{t-1}\frac{y_{i}+y_{i+1}}{2}}{\mu_{\text{ref}}\cdot(t-1)}\times 100\%$
In summary, SQ measures the quality of the final outcome, while AUC evaluates the entire discovery process by rewarding both speed and consistency. The detailed rationale for these metric designs is provided in Appendix O.2.
@@ 5.1 Performance Comparison
We use SQ to compare the capability of reaching optimal solutions, and AUC to assess discovery efficiency by measuring cumulative progress over the exploration budget.
As shown in Table 1, PiFlow achieves 49.86% AUC and 65.41% SQ on average, outperforming traditional baselines by substantial margins: +24.17% AUC and +28.70% SQ over Vanilla, and +20.58% AUC and +18.15% SQ over MPO. Against SOTA agents, PiFlow surpasses AI Researcher by 41.73% in AUC and The AI Scientist v2 by 31.72% in SQ, demonstrating superior global search capabilities across all three parameter space types (continuous, discrete, and hybrid).
As visualized in Figure 3, The AI Scientist v2 achieves the highest AUC (28.43%) in SPO due to rapid early gains, yet its trajectory stagnates, yielding a final SQ of only 29.85%—a 16.8% gap below PiFlow’s 34.85%. This illustrates entrapment in local optima. In contrast, PiFlow’s oscillatory trajectories reflect an active exploration strategy that escapes local traps. On average, PiFlow outperforms The AI Scientist v2 by 31.18% in AUC (49.86% vs. 38.01%) and 31.72% in SQ (65.41% vs. 49.66%), confirming that the observed variance stems from productive exploration rather than instability. Further details and statistical significance tests are provided in Appendix F.
Takeaway: Unlike baselines (e.g., The AI Scientist v2 (Yamada et al., 2025)), which suffer from premature stagnation, PiFlow dynamically navigates the solution space to locate high-outcome regions.
Table 2: Ablation of Test-Time Reasoning. Comparison of performance with explicit reasoning enabled (w/ Think) versus disabled (w/o Think). Bold marks the best results. Disabling the specific Think mode yields superior performance, suggesting that avoiding explicit Chain-of-Thought helps prevent cognitive fixation in this context (see discussion in Section 5.2).
Method/Setting AUC (%) SQ (%)
Qwen3-32B w/ Think 37.51 $\pm$7.70 58.76 $\pm$6.18
w/o Think 45.51 $\pm$11.19 68.86 $\pm$17.67
Qwen3-8B w/ Think 30.55 $\pm$27.45 54.49 $\pm$22.25
w/o Think 42.09 $\pm$16.55 61.59 $\pm$16.43
@@ 5.2 Ablation Study
We conduct ablations on the NHO task to evaluate two aspects: (1) the Plug-and-Play effectiveness of PiFlow, and (2) the impact of LLM reasoning modes.
[…]
Test-Time Reasoning Effect. We ablate the explicit Think mode (Chain-of-Thought) on Qwen3 models (Table 2). Counterintuitively, disabling Think yields superior results: +21.3% AUC for Qwen3-32B (37.51% vs. 45.51%) and +37.8% AUC for Qwen3-8B (30.55% vs. 42.09%). We attribute this to cognitive fixation, where explicit CoT induces path dependence, causing premature commitment to flawed assumptions. Direct inference under PiFlow’s guidance enables more flexible hypothesis generation.
Takeaway: PiFlow is a robust Plug-and-Play module that boosts performance across LLM backbones. Disabling explicit CoT reasoning synergizes with PiFlow’s strategic guidance, avoiding cognitive fixation.
Table 3: Ablation Study on the PiFlow Layer. Performance comparison with (w/) and without (w/o) the PiFlow module. Bold indicates the best performance in each setting.
@@
Figure 4: Empirical validation of theoretical alignment. (a) Average regret over time (log-log scale) closely adheres to the theoretical $\mathcal{O}(T^{-1/2})$ sublinear decay bound. (b) The positive correlation between regret and information gain confirms the dynamic shift from high-uncertainty exploration to low-regret exploitation. (c) The projected trajectory on the NHO landscape demonstrates this strategic behavior, where PiFlow traverses low-value regions to escape local optima before converging to the global peak.
5.3 Further Analyses of Robustness
To rigorously validate the proposed paradigm, we conducted an extensive suite of eight experiments analyzing PiFlow’s robustness and utility. These evaluations are structured to substantiate our three core contributions: (1) Algorithmic Robustness: We verify that the Min-Max optimization ensures resilience against epistemic uncertainty (e.g., poor initialization) and outperforms numerical baselines in ill-defined spaces; (2) Universal Applicability: We demonstrate PiFlow’s minimal-modification Plug-and-Play capability across different agent architectures and LLM backbones; and (3) Operational Efficiency: We quantify its manageable computational complexity and superior cost-effectiveness. Key findings are highlighted below:
➊ Resilience to Poor Principles Initialization. PiFlow demonstrates strong error recovery capabilities. Even when initialized with deliberately misleading principles (starting SQ below $20\%$), the system successfully identifies and discards flawed priors within $\sim 10$ iterations, recovering to converge on high-quality solutions ($\sim 70\%$ SQ), matching the performance of expert-guided runs (see Appendix M).
➋ Superiority over Bayesian Optimization. On complex tasks with ill-defined search spaces (MBO), PiFlow substantially outperforms Bayesian Optimization (BO), achieving a SQ of 84.55% compared to BO’s 38.15%, a margin of $46.4\%$, without requiring the manual parameter space engineering mandated by BO (see Appendix G).
➌ Seamless Plug-and-Play Integration. We validate the modularity of PiFlow by integrating it with ChemToolAgent (Yu et al., 2024). Without any architectural modifications, PiFlow successfully steered the agent to discover a high-potency molecule ($90.77\%$) in just 8 iterations, demonstrating its lightweight integration (see Appendix H).
➍ Negligible Computational Overhead. Theoretical analysis confirms a decision complexity of $\mathcal{O}(t^{2}\cdot d)$. Empirical profiling on NHO reveals that the PiFlow module accounts for only 9.4% of the total system runtime, efficiently delivering a 5.6x speedup in time-to-solution compared to the Vanilla baseline (see Appendix I).
➎ High Cost-Effectiveness. Compared to Vanilla MAS, PiFlow-MAS achieves a 27% reduction in total token consumption while significantly boosting performance. Crucially, the strategic overhead is minimal, with the PiFlow module itself consuming only $\sim$1.5% of the total token budget (see Appendix J).
➏ Ablations of LLMs and $\lambda$. PiFlow proves robust across diverse LLMs (e.g., Gemini-2.5-Pro (Comanici et al., 2025)), converting base reasoning capability into improved optimization outcomes. Furthermore, its behavior is predictably tunable via the trade-off parameter $\lambda$ (see Appendix L & N).
➐ Ablation of Internal Dynamics in PiFlow. As shown in Figure 9, PiFlow dynamically modulates the importance of scientific principles. It effectively discards diminishing avenues while elevating under-explored principles that reveal high information gain later, ensuring an adaptive exploration-exploitation balance (see Appendix K).
5.4 Theoretical-Empirical Alignment
The Min-Max optimization in Eq. 1 provides two theoretical guarantees (Theorem 3.2): (1) sublinear regret decay of $\mathcal{O}(\sqrt{T})$, and (2) a positive correlation between information gain and regret. We empirically validate both properties on the NHO task, as visualized in Figure 4.
(Alignment 1) Sublinear Regret Decay. The first term of Eq. 1, $\sum_{t=1}^{T}(v^{*}-f^{*}(h_{t}))$, represents cumulative regret. Theorem 3.2 guarantees that the average regret decays as $\mathcal{O}(T^{-1/2})$. Figure 4 validates this on a log-log scale: fitting yields $c\cdot T^{-0.5}$ with $c=1.37$ and $r^{2}=0.96$, confirming strong alignment. Notably, Run ID=1 exhibits a transient regret increase (iter 8–12) before resuming decay—this reflects the $\max_{f^{*}}$ adversarial term actively escaping local optima, a hallmark of the Min-Max strategy’s robustness.
(Alignment 2) Information Gain–Regret Correlation. The second term of Eq. 1, $\lambda\cdot I(h_{t};f^{*}|H_{t-1})$, maximizes mutual information for exploration. Theorem 3.2 predicts that as information gain decreases (uncertainty resolved), regret also decreases (exploitation improves). Figure 4 confirms this positive correlation: early iterations cluster in the high-gain/high-regret region (exploration phase), while later iterations concentrate in the low-gain/low-regret region (exploitation phase), demonstrating the dynamic transition governed by the trade-off parameter $\lambda$.
(Alignment 3) Trajectory Dynamics. Figure 4 visualizes the agent’s path on the PCA-projected NHO landscape. The trajectory exhibits three distinct phases that directly correspond to PiFlow’s strategic actions (Section 3.3): (i) Explore (iter 1–16): broad sampling to maximize $I(h_{t};f^{*}|H_{t-1})$; (ii) Escape (iter 16–21): traversing a low-value “valley” to avoid local optima, reflecting the $\max_{f^{*}}$ adversarial robustness; and (iii) Exploit (iter 21–24): convergence to the global peak, minimizing regret. This empirical behavior demonstrates precise alignment with the theoretical Min-Max framework.
Takeaway: Empirical analysis across all three metrics, regret decay ($r^{2}=0.96$), information-regret correlation, and trajectory dynamics, corroborates the theoretical guarantees of Theorem 3.2, validating that PiFlow’s Min-Max optimization operates as designed.
@@ 6 Conclusion
We introduce PiFlow, a Plug-and-Play framework designed to overcome the inefficiencies of unguided scientific exploration. Underpinned by a Min-Max optimization with theoretical regret guarantees, PiFlow strategically balances the exploitation of high-potential scientific principles with the exploration of epistemic uncertainty. Extensive evaluations confirm its capacity to navigate complex hypothesis spaces, significantly outperforming baselines and establishing a robust, theoretically grounded paradigm for autonomous scientific discovery.
Impact Statement
This paper presents work whose goal is to advance the field of autonomous scientific discovery. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Detailed Methodological Review
@@
A hallmark and fundamental limitation of many contemporary LLM-based agent systems in scientific discovery is their limited generalizability. As detailed in Table 4, these frameworks are often characterized by a tight coupling between their core logic and a specific scientific domain. This means their implementations, tool integrations, and most critically, their prompt engineering strategies are meticulously tailored for a single scenario, such as organic chemistry or materials science. Although systems like The AI Scientist (Lu et al., 2024) and Agent Laboratory (Schmidgall et al., 2025) demonstrate strong capabilities in scientific research, they are specifically designed for AI domain, along with a focus on the whole workflow rather than strategic decision-making problem. Consequently, transferring these systems to a new domain necessitates significant re-engineering, restricting their out-of-the-box applicability and hindering the development of truly universal systems.
Table 4: Comparison of Domain Coupling and Generalizability. Unlike domain-specific frameworks that tightly couple implementation with domain logic (e.g., specific to drug discovery), PiFlow adopts a decoupled architecture, enabling universal applicability across diverse scientific fields without re-engineering.
@@
(Ghafarollahi and Buehler, 2025) Automating alloy design and discovery using physics-aware multimodal multi-agent AI Alloy design and discovery
@@
MOOSE-Chem (Yang et al., 2024b) Utilizing Large Language Models for rediscovering unseen chemistry scientific hypotheses. Chemistry hypothesis discovery
@@
ProtAgents (Ghafarollahi and Buehler, 2024b) Protein discovery via large language model multi-agent collaborations combining physics and machine learning Protein discovery
@@
Appendix B Distinction from Advanced Prompt Engineering
We define a scientific principle following Definition 3.1. Formally, each principle is represented as a structured text proposition that can be algorithmically scored by our MinMax optimization (as exampled in Figure 1). Unlike arbitrary text strings, these principles are supposed to be with logical consistency, and this structural requirement is what distinguishes them from general prompts. In short, principles represent foundational concepts or established patterns within a domain (e.g., “higher hydrophobicity often correlates with better cell membrane penetration”). These “principles”, whether has been validated or not, can be proposed by experts or, crucially, extracted from the LLM’s own vast pre-trained knowledge. They serve as high-level starting points to generate specific, testable hypotheses.